OpenStudy (anonymous):

2x+4 if x<-1 X^2+1 if -1≤x<2 F(x) 3-x if 24 help me to find the value of x where the function f(x) is continuous.

OpenStudy (anonymous):

Does the question ask for values where f(x) is not continuous? I ask because there are an infinite number of values for x at which f(x) is continuous.

OpenStudy (anonymous):

no..it just ask where the function is continuous...

OpenStudy (anonymous):

Maybe they are limiting the question to the 3 values (1,2 and 4) that are at the boundaries of the subdomains. In that case, what you could do is draw a graph of the function and see at which of these three points the graph is continous.

OpenStudy (anonymous):

how to draw the graph? can u explain to me?

OpenStudy (anonymous):

(Correction: the subdomain boundaries are -1, 2 and 4.) Draw the line y=2x+4 to the left of x=-1. Draw the parabola \[y=x^2+1\] for x between -1 and 2. Draw the line y=3-x for x between 2 and 4. Draw a point at y=3 for x=4. Draw line y=7-2x to the right of x=4.

OpenStudy (anonymous):

ok can u explain me the feature of continuous graph?

OpenStudy (anonymous):

A function that's continuous at a particular point usually looks "connected" at that point when its graph is drawn.

OpenStudy (anonymous):

ok i will try..btw thank u for help me.. =)

OpenStudy (anonymous):

Keep in mind that even if two pieces of the function look like they come together at some value of x, the function is not be continuous if just one point is not connected there.